A proposed new venture will cost $85,000 and should produce annual cash flows of $30,000, $55,000, $40,000, and $40,000 for Years 1 to 4, respectively. The discount rate is 10 percent. What is the discounted payback period? What is IRR?
Answer #1

(1)

Discounted payback period (DPBP) is the year when the project's cumulative discounted cash flow is zero.

Year Cash Flow ($) PV Factor @10% Discounted Cash Flow ($) Cumulative Discounted Cash Flow ($)
0 -85,000 1.0000 -85,000 -85,000
1 30,000 0.9091 27,273 -57,727
2 55,000 0.8264 45,455 -12,273
3 40,000 0.7513 30,053 17,780
4 40,000 0.6830 27,321 45,100

DPBP lies between years 2 and 3.

DPBP = 2 + (Absolute value of cumulative discounted cash flow, year 2 / Discounted cash flow, year 3)

= 2 + (12,273 / 30,053)

= 2 + 0.41

= 2.41 years

(2)

IRR is computed using Excel IRR function as follows.

Year Cash Flow ($)
0 -85,000
1 30,000
2 55,000
3 40,000
4 40,000
IRR = 31.90%